OpenStudy (anonymous):

Solve by completing the square. Round to the nearest hundredth if necessary. 3x2 – 6x – 24 = 0 2.65, 3 –30, 36 4, –2 7.58, –1.58

OpenStudy (anonymous):

3(x^2-2x-8) = 0 3(x-4)(x+2)=0

OpenStudy (anonymous):

x=4, -2

OpenStudy (anonymous):

I would suggest that you search "complete the square" on youtube. There are some good videos.

OpenStudy (anonymous):

thanks i will